OpenStudy (anonymous):

Tammy’s Diner donated 6% of their weekend profit to charity. If the weekend profit of the Diner was $5,268, how much money was donated to charity?

OpenStudy (anonymous):

Money donated to charity = profit times donation percentage. Therefore, $ = $5,268 * 6 % 6% = 6/100 = .06 Therefore $ = $5,268 * .06

OpenStudy (anonymous):

$316 donated

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Looks good. YW.

OpenStudy (anonymous):

Tammy’s Diner donated 6% of their weekend profit to charity. If the weekend profit of the Diner was $5,268, how much money was donated to charity?